# Insect Identification in the Wild: The AMI Dataset

### v1

@@ 7 Conclusion
Several limitations of this work should be noted. First, while annotating AMI-Traps already occupied a considerable time for expert entomologists, further expanding it to a greater number of species and geographic regions would be helpful in further stress-testing field deployments of fine-grained classification algorithms. Second, while we focus on camera traps for nocturnal insects, it would be useful also to consider diurnal insect camera traps, for example those aimed at pollinators, which pose some additional challenges, such as multiple perspectives and a diversity of backgrounds. We hope furthermore that future works will consider more complex domain adaptation approaches building on the baselines we present here.
@@ Appendix 0.A Curating the AMI-GBIF Dataset
To ensure the quality control of the AMI-GBIF dataset, we remove images affected by issues identified as problematic for training our models: corrupted files, duplicate URLs, images from specific sources, and thumbnails. As detailed below, we also remove images containing non-adult individuals. Figure 0.A.5 shows some examples of pictures removed during our cleaning steps. Table 0.A.5 lists the sources we found containing a high percentage of unsuitable images for training models. Though species are selected based on their presence in the NE-America, C-America, and W-Europe regions, we include all GBIF observations of these species from anywhere on Earth, without any geographic filter. Figure 0.A.6 shows the distribution around the globe of the observations used for the training set of AMI-GBIF.
@@
Figure 0.A.5: Examples of pictures removed during the dataset cleaning: (a) Contains multiple specimens referenced by multiple observations, (b) a placeholder for more than 100,000 observations, (c) an unsuitable picture containing only a moth body part, (d) a thumbnail, and (e) an animal in the larval life stage.
Figure 0.A.6: Distribution of geographic locations for images in AMI-GBIF. Since we consider species from the NE-America, C-America, and W-Europe regions, observations are concentrated in these regions. However, some observations fall elsewhere, reflecting the fact that some species present in a region of interest are also found elsewhere; all global observations of that species are drawn upon in creating AMI-GBIF.
@@ Appendix 0.C Dataset
Figure 0.C.7 shows the long-tailed nature of the AMI dataset.
Figure 0.C.7: Long-tailed distribution of the data in the AMI-GBIF (left) and AMI-Traps (right) datasets. The data is skewed both at the species level and also at the genus and family levels.
@@ Appendix 0.D Data Augmentation
Figure 0.D.8 shows the data augmentation workflow in our model training pipeline. The MixRes data augmentation approach helps models generalize from AMI-GBIF training images (which are typically high resolution) to AMI-Traps test images (which are typically very low resolution).
Figure 0.D.8: The original AMI-GBIF images are stored in the WebDataset format with the smallest dimension of 450px and the other being variable to preserve the aspect ratio. First, with a probability of 0.25 each, images are downsized to a side length of 64 or 32 pixels, while the rest maintain their original size. Next, additional data augmentation techniques are applied, which include random resize crop, random horizontal flip, and RandAugment [24]. Finally, the images are resized to 128 pixels before being used as training data.
@@ Appendix 0.E Additional Experiments
Table 0.E.8: Long-tailed accuracy results for AMI-GBIF. ConvNeXt-B achieved the best performance for all regions, including a considerably higher accuracy for the medium- and few-shot species.

### v2

@@ 7 Conclusion
Several limitations of this work should be noted. First, while annotating AMI-Traps already occupied a considerable time for expert entomologists, further expanding it to a greater number of species and geographic regions would be helpful in further stress-testing field deployments of fine-grained classification algorithms. Second, while we focus on camera traps for nocturnal insects, it would be useful also to consider diurnal insect camera traps, for example those aimed at pollinators, which pose some additional challenges, such as multiple perspectives and a diversity of backgrounds. We hope furthermore that future works will consider more sophisticated domain adaptation approaches building on the baselines we present here.
@@
Acknowledgements
We thank Annie-Shan Morin, Anna Viklund, and Jamie Alison for their helpful input and support to the project.
This research was generously funded by Mila - Quebec AI Institute and Ministère de l’Économie, de l’Innovation et de l’Énergie. The work was additionally supported in part by the Canada CIFAR AI Chairs program (NSERC 585136), the Global Climate Center on AI and Biodiversity Change (ABC) (NSF OISE-2330423), the Montreal Insectarium, the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES-PROEX) - Finance Code 001, the Amazonas State Research Support Foundation - FAPEAM - through the POSGRAD project 2024/2025, the AMBER project funded by the abrdn Charitable Foundation, the Aage V. Jensen Naturfond grant no. N2024-0013, the European Union’s Horizon Europe Research and Innovation programme under grant agreement no. 101060639 (MAMBO), and a Discovery Grant by NSERC. The computing resources were provided by Mila - Quebec AI Institute, the Digital Research Alliance of Canada, and NVIDIA Corporation.
A.J., F.C., and M.J.B. led dataset creation and experiments. J.S.C. and L.P. contributed to machine learning pipeline development. N.P., F.H, J.R., M.B., M.S., J.F., A.A., Y.L., E.N., F.P.P, A.C.Z., J.A.R.S., and J.G. performed insect identification for the AMI-Traps dataset. T.A, K.B., A.G.S., M.B., Y.B., K.P.M, D. Roy, T.T.H., M.L., and D. Rolnick guided and advised the project.
@@ Appendix 0.A Curating the AMI-GBIF Dataset
To ensure the quality control of the AMI-GBIF dataset, we remove images affected by issues identified as problematic for training our models: corrupted files, duplicate URLs, images from specific sources, and thumbnails. As detailed below, we also remove images containing non-adult individuals. Figure 0.A.4 shows some examples of pictures removed during our cleaning steps. Table 0.A.5 lists the sources we found containing a high percentage of unsuitable images for training models. Though species are selected based on their presence in the NE-America, C-America, and W-Europe regions, we include all GBIF observations of these species from anywhere on Earth, without any geographic filter. Figure 0.A.5 shows the distribution around the globe of the observations used for the training set of AMI-GBIF.
@@
Figure 0.A.4: Examples of pictures removed during the dataset cleaning: (a) Contains multiple specimens referenced by multiple observations, (b) a placeholder for more than 100,000 observations, (c) an unsuitable picture containing only a moth body part, (d) a thumbnail, and (e) an animal in the larval life stage.
Figure 0.A.5: Distribution of geographic locations for images in AMI-GBIF. Since we consider species from the NE-America, C-America, and W-Europe regions, observations are concentrated in these regions. However, some observations fall elsewhere, reflecting the fact that some species present in a region of interest are also found elsewhere; all global observations of that species are drawn upon in creating AMI-GBIF.
@@ Appendix 0.C Dataset
Figure 0.C.6 shows the long-tailed nature of the AMI dataset.
Figure 0.C.6: Long-tailed distribution of the data in the AMI-GBIF (left) and AMI-Traps (right) datasets. The data is skewed both at the species level and also at the genus and family levels.
@@ Appendix 0.D Data Augmentation
Figure 0.D.7 shows the data augmentation workflow in our model training pipeline. The MixRes data augmentation approach helps models generalize from AMI-GBIF training images (which are typically high resolution) to AMI-Traps test images (which are typically very low resolution).
Figure 0.D.7: The original AMI-GBIF images are stored in the WebDataset format with the smallest dimension of 450px and the other being variable to preserve the aspect ratio. First, with a probability of 0.25 each, images are downsized to a side length of 64 or 32 pixels, while the rest maintain their original size. Next, additional data augmentation techniques are applied, which include random resize crop, random horizontal flip, and RandAugment [24]. Finally, the images are resized to 128 pixels before being used as training data.
@@ Appendix 0.E Additional Experiments
Pre-Training Ablation. Table 0.E.9 shows an ablation study to check the influence of pre-training on fine-grained classification accuracy. It is evident that using ImageNet weights and pre-training on all AMI-GBIF data followed by region-specific fine-tuning achieves the best performance. Training from scratch also shows good results but requires much longer epochs.
Analysis of Accuracy v/s Confidence: One important point for practical use cases is how accuracy changes if predictions are only accepted above a certain "confidence threshold". Thresholding predictions can increase the reliability of model outputs for downstream applications, though it comes with a trade-off of rejecting a certain fraction of predictions. Figure 0.E.8 shows the curves of the accuracy and rejection rate as functions of the confidence threshold.
@@
Table 0.E.8: Micro-averaged long-tailed accuracy results for AMI-GBIF. ConvNeXt-B achieved the best performance for all regions, including a considerably higher accuracy for the medium- and few-shot species.
@@
Table 0.E.9: Ablation study results for different choices of pre-training. PT and FT stands for pre-training and fine-tuning respectively. When training from scratch, we use a schedule of 90 epochs and an initial learning rate of 0.01. As the differences in results may be due to the number of epochs, we keep the total epochs consistent (30 or 90) when pretraining on all AMI-GBIF data and then fine-tuning on regional checklists. The model architecture is ResNet50 with ImageNet weights loaded from the timm library, wherever applicable. The results are species accuracy for the NE-America region.
[…]
60 30 90.28$\pm$0.78 75.06$\pm$0.42 67.34$\pm$0.71
20 10 90.60$\pm$0.05 73.57$\pm$0.52 66.16$\pm$0.51
Figure 0.E.8: The left y-axis compares the top-1 accuracy as a function of model (ResNet50) confidence, while the right y-axis shows the rejection rate (ratio of rejected samples to total samples) for the corresponding thresholds for the NE-America (left) and W-Europe regions (right). At a confidence of 50%, the model has over 90% accuracy at all taxonomic levels. This analysis is skipped for the C-America which has sparse taxonomic coverage in the AMI-T dataset.